# smallequation

### v10

@@ 2.1 Structural Causal Model
Here $[d]=\{0,1,...,d\}$. We use $X_{i}$ to represent the $i$th variable as well as the $i$th node in $\mathcal{G}$, and $\mathbf{pa}(i)$ stands for the parents of $i$, and $\mathbf{ch}(i)$ is the children of node $X_{i}$ regarding $\mathcal{G}$. In addition, we have $U_{i}\perp\!\!\!\perp X_{i}$ and $U_{i}\perp\!\!\!\perp U_{j},\forall i\neq j$. $f_{i}$ is a mapping from $\mathbb{R}^{|\mathbf{pa}(i)|+1}\rightarrow\mathbb{R}$. $\mathcal{X}_{i}$, $\mathbf{\mathcal{Z}}_{i}$, and $\mathcal{U}_{i}$ stand for the value domain of $X_{i}$, $\mathbf{Z}_{i}$, and $U_{i}$, respectively. We further assume that the expectation of $X_{i},i\in[d]$, i.e. $\mathbb{E}[X_{i}]$, exists. Existing CBO methods [1; 22] usually assume that the exogenous variable is an ANM [11] to the data generation mechanism, i.e., $X_{i}=f_{i}(\mathbf{Z}_{i})+U_{i},\ \text{with}\ U_{i}\sim\mathcal{N}(0,1).$
@@ 2.2 Intervention
where $A_{i}$ is the continuous action variable for $X_{i}$ if $X_{i}\in\mathbf{I}$. Action $A_{i}$ with the corresponding value domain $\mathcal{A}_{i}$ controls the change of the data generation mechanism of node $X_{i}$. The soft intervention can be represented with Pearl’s do-operation [18] $do\big{(}\mathbf{X}_{\mathbf{I}}:=\mathbf{f}(\mathbf{Z}_{\mathbf{I}},A,U_{\mathbf{I}})\big{)}$
@@ 4.1 Exogenous Variable Recovery for One Node
Figure 2: Structure in one $\tau$-SCM node. $\mathbf{Z}$ is the parent set of $X$, and our algorithm learns an encoder $h$ and a decoder $g$ so that the surrogate $\widehat{U}=h(\mathbf{Z},X)$ and $X=g(\mathbf{Z},\widehat{U})$.
@@
(EDS) Let $(\mathbf{Z},U,X,f)$ be a $\tau$-SCM. Let $\phi()$ be a regression model, $\phi():\mathbf{\mathcal{Z}}\to\mathcal{X}$. We assume that $\mathbb{E}[X]$ exists and $\phi()$ has sufficient expressiveness to capture the regression mean $\mu_{\phi}()$ and variance $\sigma_{\phi}()$. We define $(\widehat{U},\phi,h,g)$ as encoder-decoder surrogate (EDS) model of the exogenous variable $U$ with the encoder $h():\mathcal{Z}\times\mathcal{X}\to\widehat{\mathcal{U}}$, $\widehat{U}:=h(\mathbf{Z},X):=\frac{X-\mu_{\phi}(\mathbf{Z})}{\sigma_{\phi}(\mathbf{Z})}$; and the decoder $g():\mathcal{Z}\times\widehat{\mathcal{U}}\to\mathcal{X}$, $X:=g(\mathbf{Z},\widehat{U})$.
With the observations of $X$ and its parent $\mathbf{Z}$, our method learns the encoder $h()$ that aims to recover the value of $U$, i.e., $\widehat{u}=h(\mathbf{z},x)$; our algorithm also learns the decoder $x=g(\mathbf{z},\widehat{u})$ that provides us a surrogate model of the mechanism $f()$. Theorem 4.1 shows that the surrogate values of the exogenous variable $U$ can be recovered from the observations of node $X$ under the DGM assumption of $f$.
@@
We use the distribution of $\widehat{U}=\mathbf{s}(U)=h(\mathbf{Z},X)$, i.e. $p(\widehat{U})$, to represent $p(U)$ in the surrogate model. In our surrogate model, $f()$ is approximated with the learned $g()$ and $\widehat{u}$, i.e., $x=f(\mathbf{z},u)=g(\mathbf{z},\widehat{u})=g(\mathbf{z},\mathbf{s}(u))$. Figure 3 gives the relationship between different mechanisms. These features substantially extend the identification of exogenous variable $U$ to much broader nonlinear and non-monotonic data-generation mechanism classes compared to ANM (linear) and BGM (monotonic, [16]). DGM is significantly different from the ANM [11] mechanism given by $X=f(\mathbf{pa}(X))+U$, which is a linear monotonic function of $U$ and widely used by many existing BO and CBO methods. We use EDS to represent the $\tau$-SCMs with the exogenous distribution that can be approximately recovered by using EDS (with or without the condition $\widehat{U}\perp\!\!\!\perp\mathbf{Z}$) and Figure 3 gives the scopes of different mechanisms including EDS.
@@ 4.2 Implementation of Exogenous Distribution Learning
We use a Gaussian Mixture model to fit the recovered distribution of $\widehat{U}$, i.e. $p(\widehat{U})$. We use $p(\widehat{U})$ to replace $p(U)$ in the computing of the probabilistic surrogate objective. For the nodes in SCM $\mathcal{M}$, let $\mathbf{G}=\{g_{i}\}_{i=0}^{d}$, and $\mathbf{H}=\{h_{i}\}_{i=0}^{d}$.
@@ 5.2 Acquisition Function
where the expectation is taken over $p(\widehat{U})$. In (8), $\beta_{t}$ controls the tradeoff between exploration and exploitation of Algorithm 1. The UCB-based algorithm is a classic strategy that is widely used in BO and stochastic bandits [14; 20]. The proposed algorithm adapts the “optimism in the face of uncertainty” (OFU) strategy by taking the expectation of the UCB as part of the acquisition process.
@@ Appendix A Analysis on Exogenous Variable Recovery
Let $(X,\mathbf{Z},U,f)$ be a $\tau$-SCM. Let $\rho():\mathcal{X}\times\mathcal{Z}\to\mathbb{R}^{1}$ be a predefined function regarding $X$ and $\mathbf{Z}$, and $\phi()$ be a regression model with $\phi():\mathcal{Z}\to\rho(\mathcal{X},\mathcal{Z})$. We define an encoder function $h():\mathcal{Z}\times\mathcal{X}\to\widehat{\mathcal{U}}$ with $\widehat{U}:=h(\mathbf{Z},X):=\rho(X,\mathbf{Z})-\phi(\mathbf{Z})$. The decoder is $g():\mathcal{Z}\times\widehat{\mathcal{U}}\to\mathcal{X}$, i.e., $X=g(\mathbf{Z},\widehat{U})$. Let $\rho()$ maps the values of $X$ and $\mathbf{Z}$ to an additive function of $\mathbf{Z}$ and $U$, i.e., $\rho(X,\mathbf{Z})=\rho_{1}(\mathbf{Z})+\rho_{2}(U)$. Then $\widehat{U}=h(\mathbf{Z},X)=\rho_{2}(U)-\mathbb{E}[\rho_{2}(U)]$, and $\widehat{U}\perp\!\!\!\perp\mathbf{Z}$.
@@
[…]
@@
For an ANM [11] model $X=f(\mathbf{Z})+U$, we have $\rho(X,\mathbf{Z})=X$, $\rho_{1}(\mathbf{Z})=f(\mathbf{Z})$, and $\rho_{2}(U)=U$, then $\widehat{U}=h(\mathbf{Z},X)=U-\bar{U}$
@@
For a model $X=2Ze^{-U}-e^{-Z}$, we have $\rho(X,Z)=\log(X+e^{-Z})$, $\rho_{1}(Z)=\log(2Z)$, and $\rho_{2}(U)=-U$, then $\widehat{U}=h(Z,X)=-U+\bar{U}$
@@
[…]
@@ Appendix B Regret Analysis
Our analysis follows the study in [22]. In the DAG $\mathcal{G}$ over $\{X_{i}\}_{0}^{d}$, let $N$ be the maximum distance from a root to $X_{d}$, i.e., $N=\max_{i}\text{dist}(X_{i},X_{d})$. Here $\text{dist}(\cdot,\cdot)$ is a measure of the edges in the longest path from $X_{i}$ to the reword node $X_{d}$. Let $M$ denote the maximum number of parents of any variables in $\mathcal{G},M=\max_{i}|\mathbf{pa}(i)|$. Let $L_{t}$ be a function of $L_{g}$, $L_{\sigma_{g}}$. According to Theorem 4.1, with the EDS structure given in Figure 2 in the main text, the exogenous variable and its distribution can be recovered. For each observation of the dynamic surrogate model, we assume the sampling of $p(\widehat{U})$, $\tilde{\widehat{u}}=\mathbf{s}(\tilde{u})=\mathbf{s}(u)$.

### v11

# Causal Bayesian Optimization via Exogenous Distribution Learning

@@ 2.1 Structural Causal Model
Here $d]=\{0,1,...,d\}$. We use $X_{i}$ to represent the $i$th variable as well as the $i$th node in $\mathcal{G}$, and $\mathbf{pa}(i)$ stands for the parents of $i$, and $\mathbf{ch}(i)$ is the children of node $X_{i}$ regarding $\mathcal{G}$. In addition, we have $U_{i}\perp\!\!\!\perp\mathbf{Z}_{i}$ and $U_{i}\perp\!\!\!\perp U_{j},\forall i\neq j$. $f_{i}$ is a mapping from $\mathbb{R}^{|\mathbf{pa}(i)|+1}\rightarrow\mathbb{R}$. $\mathcal{X}_{i}$, $\mathbf{\mathcal{Z}}_{i}$, and $\mathcal{U}_{i}$ stand for the value domain of $X_{i}$, $\mathbf{Z}_{i}$, and $U_{i}$, respectively. We further assume that the expectation of $X_{i},i\in[d]$, i.e. $\mathbb{E}[X_{i}]$, exists. Existing CBO methods [1; 22] usually assume that the exogenous variable is an ANM [11] to the data generation mechanism, i.e., $X_{i}=f_{i}(\mathbf{Z}_{i})+U_{i},\ \text{with}\ U_{i}\sim\mathcal{N}(0,1).$
@@ 2.2 Intervention
where $A_{i}$ is the continuous action variable for $X_{i}$ if $X_{i}\in\mathbf{I}$. Action $A_{i}$ with the corresponding value domain $\mathcal{A}_{i}$ controls the change of the data generation mechanism of node $X_{i}$. The soft intervention can be represented with Pearl’s do-operation [18] $do\big(\mathbf{X}_{\mathbf{I}}:=\mathbf{f}(\mathbf{Z}_{\mathbf{I}},A,U_{\mathbf{I}})\big)$ .
@@ 4.1 Exogenous Variable Recovery for One Node
[…]
@@
[…]
@@
We use the distribution of $\widehat{U}=\mathbf{s}(U)=h(\mathbf{Z},X)$, i.e. p(U^)p($\widehat${U}), to represent $p(U)$ in the surrogate model. In our surrogate model, $f()$ is approximated with the learned $g()$ and $\widehat{u}$, i.e., $x=f(\mathbf{z},u)=g(\mathbf{z},\widehat{u})=g(\mathbf{z},\mathbf{s}(u))$. Figure 3 gives the relationship between different mechanisms. These features substantially extend the identification of exogenous variable $U$ to much broader nonlinear and non-monotonic data-generation mechanism classes compared to ANM (linear) and BGM (monotonic, [16]). DGM is significantly different from the ANM [11] mechanism given by $X=f(\mathbf{pa}(X))+U$, which is a linear monotonic function of $U$ and widely used by many existing BO and CBO methods. We use EDS to represent the $\tau$-SCMs with the exogenous distribution that can be approximately recovered by using EDS (with or without the condition $\widehat{U}\perp\!\!\!\perp\mathbf{Z}$) and Figure 3 gives the scopes of different mechanisms including EDS.
@@ 4.2 Implementation of Exogenous Distribution Learning
We use a Gaussian Mixture model to fit the recovered distribution of $\widehat{U}$, i.e. p(U^)p($\widehat${U}). We use p(U^)p($\widehat${U}) to replace $p(U)$ in the computing of the probabilistic surrogate objective. For the nodes in SCM $\mathcal{M}$, let $\mathbf{G}=\{g_{i}\}_{i=0}^{d}$, and $\mathbf{H}=\{h_{i}\}_{i=0}^{d}$.
@@ 5.2 Acquisition Function
where the expectation is taken over p(U^)p($\widehat${U}). In (6), $\beta_{t}$ controls the tradeoff between exploration and exploitation of Algorithm 1. The UCB-based algorithm is a classic strategy that is widely used in BO and stochastic bandits [14; 20]. The proposed algorithm adapts the “optimism in the face of uncertainty” (OFU) strategy by taking the expectation of the UCB as part of the acquisition process.
@@ Appendix A Analysis on Exogenous Variable Recovery
[…]
@@
$\phi(\mathbf{z})$ $=\mathbb{E}[\rho(X,\mathbf{z})]=\mathbb{E}[\rho_{1}(\mathbf{z})+\rho_{2}(U)]=\int\big(\rho_{1}(\mathbf{z})+\rho_{2}(u)\big)p(u)du$
@@
For an ANM [11] model $X=f(\mathbf{Z})+U$, we have $\rho(X,\mathbf{Z})=X$, $\rho_{1}(\mathbf{Z})=f(\mathbf{Z})$, and $\rho_{2}(U)=U$, then $\widehat{U}=h(\mathbf{Z},X)=U-\bar{U}$ .
@@
For a model $X=2Ze^{-U}-e^{-Z}$, we have $\rho(X,Z)=\log(X+e^{-Z})$, $\rho_{1}(Z)=\log(2Z)$, and $\rho_{2}(U)=-U$, then $\widehat{U}=h(Z,X)=-U+\bar{U}$ .
@@
$\mu_{\phi}(\mathbf{z})$ $=\mathbb{E}[X(\mathbf{z},U)]=\int\big(f_{a}(\mathbf{z})+f_{b}(\mathbf{z})f_{c}(u)\big)p(u)du$
$=f_{a}(\mathbf{z})+\int f_{b}(\mathbf{z})f_{c}(u)p(u)du$
@@ Appendix B Regret Analysis
Our analysis follows the study in [22]. In the DAG $\mathcal{G}$ over $\{X_{i}\}_{0}^{d}$, let $N$ be the maximum distance from a root to $X_{d}$, i.e., $N=\max_{i}\text{dist}(X_{i},X_{d})$. Here $\text{dist}(\cdot,\cdot)$ is a measure of the edges in the longest path from $X_{i}$ to the reword node $X_{d}$. Let $M$ denote the maximum number of parents of any variables in $\mathcal{G},M=\max_{i}|\mathbf{pa}(i)|$. Let $L_{t}$ be a function of $L_{g}$, $L_{\sigma_{g}}$. According to Theorem 4.1, with the EDS structure given in Figure 2 in the main text, the exogenous variable and its distribution can be recovered. For each observation of the dynamic surrogate model, we assume the sampling of p(U^)p($\widehat${U}), $\tilde{\widehat{u}}=\mathbf{s}(\tilde{u})=\mathbf{s}(u)$.